# Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector

### v1

@@ Abstract
The Circular Electron-Positron Collider (CEPC), a proposed next-generation $e^{+}e^{-}$ collider to enable high-precision studies of the Higgs boson and potential new physics, imposes rigorous demands on detector technologies, particularly the vertex detector. JadePix-3, a prototype Monolithic Active Pixel Sensor (MAPS) designed for the CEPC vertex detector, targets a spatial resolution of 3 $\rm{\upmu m}$ to meet the key requirement. This paper presents a detailed laboratory-based characterization of the JadePix-3 sensor, focusing on the previously under-explored effects of substrate reverse bias voltage on key performance metrics: charge collection efficiency, average cluster size, and detection efficiency. Systematic testing demonstrated that JadePix-3 operates reliably under reverse bias, exhibiting a reduced input capacitance, an expanded depletion region, enhanced charge collection efficiency, and a lower fake-hit rate. Simultaneously, an extensive methodology for MAPS characterization is established, which determined the optimal operational working conditions for the JadePix-3 sensor. Furthermore, these findings confirm the sensor’s potential for high-precision particle tracking and vertexing at the CEPC while offering valuable references for future iterational R&D of the JadePix series.
@@ 1 Introduction
The discovery of the Higgs boson has spurred interest in circular $e^{+}e^{-}$ colliders as premier facilities for high-precision studies of the Standard Model Higgs boson and other fundamental particles. Planned designs aim to cover a wide range of center-of-mass energies—from the Z pole (91 GeV) and WW threshold (161 GeV) to the Higgs factory region (240 GeV) and the $t\bar{t}$ threshold (340–365 GeV)—using an $e^{+}e^{-}$ accelerator with a circumference of approximately 100 km. In addition, the same tunnel could later accommodate a hadron collider with a center-of-mass energy nearly 100 TeV. The Circular Electron-Positron Collider (CEPC) [1, 2] has been proposed to pursue this road map, envisioned to investigate the Higgs boson further and reveal new physics potentially.
CEPC’s ambitious physics objectives impose stringent performance requirements on its detectors—particularly the vertex detector. As outlined in Table 1, the CEPC vertex detector must achieve exceptional spatial resolution, maintain a low material budget, minimize power consumption, handle high hit densities, and exhibit robust radiation tolerance to cope with the beam-induced background near the interaction point. In this context, only Monolithic Active Pixel Sensors (MAPS) offer a promising solution to meet these demanding criteria [3, 4].
Therefore, considerable R&D efforts have been undertaken to develop advanced pixel sensor prototypes with monolithic technology for the CEPC vertex detector, such as the TaichuPix series [5, 6, 7] and the JadePix series. The JadePix series is specifically developed to meet the CEPC vertex detector’s demanding criteria, offering high position resolution along with enhanced efficiency. Building on the progressive improvements of its predecessors—JadePix-1, which was to verify the sensor fabrication and optimize the pixel architecture [8, 9], and JadePix-2, which featured an improved analog front‐end design [10]—JadePix-3 represents the latest development and aims for a position resolution of 3 $\rm{\upmu m}$ [11]. The next part presents the detailed design specifications of JadePix-3.
Table 1: Demands for the CEPC vertex detector [1].
@@
NIEL, $\rm 6.2\times 10^{12}n_{eq}/(cm^{2}year)$
Figure 1: (a) Layout of the entire sensor mask; (b) Circuit topology for the analog front-end [11].
@@ 1.1 Overview of JadePix-3 design
The JadePix-3 is designed for a high spatial resolution of 3 $\rm{\upmu m}$. With a pixel matrix consisting of 512 rows and 192 columns, the sensor covers a sensitive area of 8.19 $\times$ 4.85 $\rm{mm^{2}}$ as shown in Fig. 1. The sensor is divided into four distinguishable sectors in the column direction. As presented in Tab. 2, while the row pitch is uniformly set at 16 $\rm{\upmu m}$ across all four sectors, the column pitch varies between 23.11 $\rm{\upmu m}$ in sector 2 and 26 $\rm{\upmu m}$ in sectors 0, 1, and 3. The pixel arrangement in each sector is mirrored, with shared vertical metal lines for biasing and control signals.
Fig. 1 demonstrates JadePix-3’s front-end circuit design. As described in Ref. [11] in detailed, the initial amplification stage utilizes a cascode amplifier, which provides high gain while minimizing noise, with its output referred to as OUT_A. Subsequently, the amplified signal is fed into a discriminator stage, where it is compared against a predefined threshold. The resulting digital output, designated as OUT_D, enables efficient hit detection.
@@
0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 SR flipflop
@@
2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 SR flipflop, w/o MASK
@@
The chip employs both rolling shutter (RS) and global shutter (GS) readout sequence modes. In the conventional rolling shutter mode, one row is activated for reading at a time [12]. This process requires 192 $\rm ns$ to transmit hit data and reset the pixels, and scanning the entire 512-row matrix takes 98.3 $\rm{\upmu s}$. While in global shutter, all 512 rows are frozen for a period of time during readout.
The JadePix-3’s sensor design is shown in Fig. 2. A reverse bias $V_{\rm sub}$ can be applied on P-well and P-substrate, expanding the depletion region around the N-well electrode.
@@ 1.2 Characterization Motivation
The performance of MAPS is significantly influenced by the bias voltage applied to the substrate. This bias has a multifaceted impact: on one hand, it directly increases the size of the depletion zone and reduces the sensor’s input capacitance [14, 15]. On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the pixel circuitry tends to increase due to the body effect, consequently altering the circuit’s operating status. Therefore, adjusting the circuit’s operating point after applying substrate bias and investigating the resulting impact on these key performance metrics becomes essential. These metrics are vital for the R&D of the CEPC vertex detector.
@@
This work also introduces an integrated and systematic test methodology (detailed in Section 2), which was established to guide the subsequent comprehensive, lab-based characterization of MAPS behavior under varying reverse bias conditions. While most studies hold the substrate bias fixed or explore a narrow range [16, 17], our approach systematically sweeps the reverse bias from 0 V to –6 V under identical threshold settings. This methodology establishes comparable operating conditions across different bias voltages, thereby facilitating an unambiguous assessment of performance changes predominantly attributable to the bias variation itself.
2 Workflow for Comprehensive Characterization
This work employs a structured, integrated methodology (see Fig. 3):
  • 1.
    Phase 1: Measurement Setup, which is introduced in Section 3, focuses on the preparation stage and establishes a stable experimental setup that minimizes mechanical and environmental perturbations, thereby ensuring reliable test conditions.
  • 2.
    Phase 2: Working-Point Tuning is introduced in Section 4. This phase establishes a series of optimal operating points through a two-stage procedure.
    1. (a)
      Target threshold: Target working-points are defined as distinct operating configurations, one for each applied bias level, that all share a consistent average threshold. This is achieved by applying reverse bias and meticulously tuning DAC parameters to meet the predefined target threshold.
    2. (b)
      ITHR scan: Subsequently, an ITHR scan is performed from each of these target working-points to generate an expanded set of conditions, offering a range of distinct threshold levels for each bias.
    This tuning not only confirms the feasibility of operating under varying reverse biases but also lays the groundwork for all subsequent measurements.
  • 3.
    In Phase 3: Comprehensive Performance Evaluation, which is illustrated in Section 5, three sequentially interlinked investigations were carefully designed and performed:
    1. (a)
      Sensor Electrical Properties Characterization: Detailed measurements of leakage current and input capacitance—both regarded as critical sensor performance metrics—were conducted to understand the intrinsic electrical behavior of the sensor under bias.
    2. (b)
      Noise Performance Analysis: Fake-hit rate is measured to assess noise characteristics, since maintaining a low fake-hit rate is critical for ensuring the effectiveness of hit data for the subsequent measurements.
    3. (c)
      Detection Response Evaluation: The sensor’s response is further assessed by exposing it to external sources, including radioactive sources and an infrared laser. The laser tests, in particular, provide high-resolution mapping of in-pixel charge collection efficiency and serve as a proxy for evaluating sensor performance in response to external stimuli.
Figure 3: Detailed workflow to comprehensively characterize the MAPS sensor.
This integrated workflow not only delivers a robust, reproducible characterization of JadePix-3 in the laboratory—fully capturing how reverse bias shapes every aspect of its response—but also provides a transferable framework for future MAPS R&D.
3 Experimental Setup
@@
The JadePix-3 chip has been wire-bonded to an adaptor board and then connected to an external FPGA-based DAQ system (see Fig. 4). The FPGA board configures the chip parameters and transfers data to the PC via Ethernet. A shock-absorbing platform was employed to maintain system stability and prevent displacement of the test source, either a radioactive source or a laser source, and the sensor during the measurements. A source meter is used to apply reverse bias ranging from 0 to -6 V to the substrate.
@@
Figure 4: (a) Schematic view of the test system; (b) Setup of the infrared laser test system. All devices are fixed or placed on an optical table.
@@
To proceed laser experiment under the demand of high spatial resolution, a laser test system featuring three high-precision motor controller is deployed (see Fig. 4). The spatial position in three orthogonal directions of the laser source is regulated precisely by these three motors, each has a repeatable position resolution of 1 $\rm{\upmu m}$ and a minimum step size of 0.1 $\rm{\upmu m}$. Details about the experiment parameters and test method will be discussed thoroughly in Section 5.5.
4 Working-point Selection
@@
This section details the methodology for selecting operational working-points of the JadePix-3 sensor. The primary objective is to establish uniform thresholds that enable the comparison of the sensor’s performance across varying substrate reverse biases. Central to this process is the accurate measurement of pixel thresholds, achieved using the S-curve method.
This methodology involves giving each pixel a certain number of electrical pulses and counting the hits while increasing the injected charge, which is regulated as $V_{\rm pulse}\cdot C_{\rm test}$, where $C_{\rm test}$ is the injection capacitor that has a designed value of $0.18~{}{\rm fF}$ marked in Fig. 1, and $V_{\rm pulse}$ is the external voltage for test-pulse injection.
As the injected charge approaches the pixel threshold, the hit count rises along an S-shaped curve, typically modeled by fitting an error function. The pixel threshold is then defined as the injected charge level at which this fitted error function reaches half of its maximum value. Threshold variations between pixels are quantified by the fixed pattern noise (FPN), defined as the standard deviation of the mean threshold. Furthermore, the temporal noise (TN) is characterized by the RMS value of the derivative of the S-curve, providing a measure of the sum of both the fluctuations on the injected signal and the fluctuations on the pixel threshold itself. The final threshold value is converted to electrons via the designed value of $C_{\rm test}$.
4.1 Target Working-Point Tuning
@@
To thoroughly assess the impact of reverse bias on sensor performance, it is essential to tune the sensor to target working-points that exhibit a uniform average threshold across various substrate biases. Achieving such consistency poses a challenge due to the complex interplay of multiple parameters influencing the analog front-end. Therefore, the operating conditions—particularly the DAC settings that control the front-end state—must be meticulously adjusted.
@@
This work focuses on the characterization of sectors 0-2. Sector 3 is excluded from the present discussion; its front-end current was increased for radiation tolerance evaluations, leading to significantly higher power consumption. The analog front-end design is identical for sectors 0-2, with variations existing only in their digital logic blocks for pixel functions and in their respective pixel pitches (see Tab. 2).
@@
The M6 MOSFET, biased by VCLIP, clamps excessively high OUT_A signals (see Fig. 1). Given that the amplitude of OUT_A increases significantly with reverse bias due to a reduction in sensor capacitance, VCLIP is set to 1.307 V. This is essentially its maximum value and is chosen to eliminate the clamping effect of M6 on the OUT_A amplitude during characterization. To compensate for the increase of the M8 gate threshold induced by substrate reverse bias, VCASN2 is increased from its nominal 400 mV to 722 mV, thereby restoring M8’s switching threshold to its intended value.
@@
Settings $V_{\rm sub}$ VCASN/mV ITHR/nA VCLIP/mV VCASN2/mV
@@
The threshold distributions for these target working-points are presented in Fig. 5. It is observed that these distributions at non-zero reverse biases are noticeably broader compared to zero bias, indicating increased pixel-to-pixel threshold variations. This observation is further substantiated by the FPN analysis shown in Fig. 6, where FPN increases with reverse bias and quantitatively reflects the growing differences in threshold across the pixel matrix. The corresponding TN analysis is shown in Fig. 6. An increase in TN is observed under reverse bias conditions. The underlying reasons for this behavior will be discussed in Section 5.2 along with the change of OUT_A signal amplitude.
@@
4.2 ITHR Scan
To investigate sensor performance across a range of detection thresholds, the ITHR value was systematically scanned starting from each of these target working-points. This procedure yielded a series of operating conditions—7 distinct $V_{\rm sub}$ levels, each with 11 different ITHR settings—resulting in a set of working-points with varying thresholds.
Figure 7 illustrates the impact of the ITHR value on the mean threshold, FPN and TN, respectively. Both FPN and TN generally exhibit an increasing trend with higher threshold values.
@@ 5 Experimental Results and Discussion
Reducing sensor capacitance and expanding the depletion zone are the most intrinsic and direct consequences of applying a reverse bias on the substrate. After adjusting thresholds under different reverse biases and determining a set of working-points by ITHR value scan described in Sec. 4.2, a thorough sensor characterization regarding the impact of the substrate reverse bias voltage was conducted, including leakage current, analog response, fake-hit rate and charge collection efficiency.

### v2

# Further Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector

@@ Abstract
The Circular Electron-Positron Collider (CEPC), a proposed next-generation $e^{+}e^{-}$ collider to enable high-precision studies of the Higgs boson and potential new physics, imposes rigorous demands on detector technologies, particularly the vertex detector. JadePix-3 is a prototype Monolithic Active Pixel Sensor (MAPS) designed for the CEPC vertex detector. This paper presents a detailed laboratory-based characterization of the JadePix-3 sensor, focusing on the previously under-explored effects of substrate reverse bias voltage on key performance metrics: charge collection efficiency, average cluster size, and detection efficiency. Systematic testing demonstrated that JadePix-3 operates reliably under reverse bias, exhibiting a reduced input capacitance, an expanded depletion region, enhanced charge collection efficiency, and a lower fake-hit rate. These findings confirm the sensor’s potential for high-precision particle tracking and vertexing at the CEPC while offering valuable references for future iterational R&D of the JadePix series.
@@
{highlights}
Comprehensive laboratory characterization of the JadePix-3 CMOS pixel prototype for the vertex detector of the Circular Electron-Positron Collider
Measurements performed under various reverse bias voltages from 0 to –6 V with dedicated working point optimization at each bias voltage
Substrate reverse bias significantly improves the sensor performance such as the charge collection efficiency and the fake-hit rate
@@ 1 Introduction
Following the discovery of the Higgs boson, circular $e^{+}e^{-}$ colliders have garnered significant interest as premier facilities for high-precision studies of the Standard Model and searches for new physics. The Circular Electron-Positron Collider (CEPC) [1, 2] has been proposed to advance this frontier, with an ambitious physics program centered on detailed investigations of the Higgs boson.
The ambitious physics objectives of the CEPC impose stringent performance requirements on its detector systems, particularly on the vertex detector located nearest to the interaction point. As outlined in Table 1, the vertex detector must satisfy a challenging combination of specifications: exceptional spatial resolution, a minimal material budget, low power consumption, high hit-rate capability, and robust radiation tolerance against beam-induced backgrounds. Meeting these multifaceted demands is a significant technological challenge. In this context, Monolithic Active Pixel Sensors (MAPS) stand out as a uniquely promising technology capable of fulfilling all these criteria [3, 4]. Consequently, considerable research and development efforts are underway to advance monolithic sensor prototypes for the CEPC. The JadePix sensor series, developed specifically for this purpose, is a direct result of these efforts. This paper focuses on the JadePix-3 prototype, the design specifications of which are briefly summarized below, with further details available in Ref. [5].
Table 1: Requirements for the CEPC vertex detector [1].
@@
NIEL, $\rm 6.2\times 10^{12}~\text{1~MeV}~n_{eq}/(cm^{2}\cdot year)$
@@ 1.1 Overview of JadePix-3 design
The JadePix-3 sensor, developed with TowerJazz 180 nm CMOS imaging technology, covers a pixel matrix consisting of 512 rows and 192 columns. The sensor is divided into four distinguishable sectors in the column direction. While the row pitch is uniformly set at 16 $\rm{\upmu m}$ across all four sectors, the column pitch varies between 23.11 $\rm{\upmu m}$ in sector 2 and 26 $\rm{\upmu m}$ in sectors 0, 1, and 3. The variation in the column pitch arises from the implementation of different digital logic blocks within the pixels. These blocks were designed to evaluate the performance of R-S flipflops versus D flipflops, as well as the inclusion or exclusion of a MASK configuration register, requiring adjustments to the pixel layout. All these differences are summarized in Tab. 2.
@@
0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 R-S flipflop
@@
2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 R-S flipflop, w/o MASK
@@
In the JadePix-3 sensor design, each pixel contains a P-well that allows for substrate biasing through a surface contact, as shown schematically in Fig. 1. This P-well is electrically contiguous with the underlying deep P-well and the P-type epitaxial layer. This configuration ensures that the entire P-type substrate is held at the applied potential, $V_{\rm sub}$. The pads connected to these P-wells are designed without ESD protection to allow for the application of a wide range of substrate bias voltages. When a reverse bias $V_{\rm sub}$ is applied on P-well and P-substrate, the depletion region around the N-well electrode expands.
@@ 1.2 Characterization Motivation
Figure 2: Circuit topology for the JadePix-3 analog front-end [5].
The performance of MAPS is significantly influenced by the bias voltage applied to the substrate. This bias has a multifaceted impact: on one hand, it directly increases the size of the depletion zone and reduces the sensor’s input capacitance [7]. On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the pixel circuitry (shown in Fig. 2) tends to increase due to the body effect, consequently altering the circuit’s operating status. Therefore, adjusting the circuit’s operating point after applying substrate bias and investigating the resulting impact on these key performance metrics becomes essential. These metrics are vital for the R&D of the CEPC vertex detector.
@@
This work also introduces an integrated and systematic test methodology (detailed in Section 2.2), which was applied to guide the subsequent lab-based characterization of MAPS behavior under varying reverse bias conditions. Our approach systematically sweeps the reverse bias from 0 V to –6 V under identical threshold settings. This methodology establishes comparable operating conditions across different bias voltages, thereby facilitating an unambiguous assessment of performance changes predominantly attributable to the bias variation itself.
2 Experimental Methods
2.1 Measurement Setup
@@
The experimental setup, which is based on the one described in Ref. [5], is shown in Fig. 3. The JadePix-3 chip has been wire-bonded to a carrier board and then connected to an external FPGA-based DAQ system (see Fig. 3). The FPGA board configures the chip parameters and transfers data to the PC via Ethernet. A shock-absorbing platform was employed to maintain system stability and prevent displacement of the test source, either a radioactive source or a laser source, and the sensor during the measurements. A source meter is used to apply reverse bias ranging from 0 to -6 V to the substrate. The upper limit of -6 V was chosen as a conservative value to ensure a stable operation, based on breakdown voltages reported for other sensors fabricated in the same 180 nm process [7].
@@
Figure 3: (a) Schematic view of the test system; (b) Setup of the infrared laser test system. There is a window in the PCB to allow radiative source or laser beam incident to the JadePix-3 sensor. All devices are fixed or placed on an optical table.
@@
To conduct laser experiment under the demand of high spatial resolution, a laser test system featuring three high-precision motor controllers is deployed (see Fig. 3). The spatial position in three orthogonal directions of the laser source is regulated precisely by these three motors, each has a repeatable position resolution of 1 $\rm{\upmu m}$ and a minimum step size of 0.1 $\rm{\upmu m}$. Details about the experiment parameters and test method will be discussed in Section 3.5.
2.2 Working-point Selection
@@
This part details the methodology for selecting operational working-points of the JadePix-3 sensor. The primary objective is to establish uniform thresholds that enable the comparison of the sensor’s performance across varying substrate reverse biases. Central to this process is the accurate measurement of pixel thresholds, achieved using the S-curve method.
The operational parameters of the pixels—threshold, fixed pattern noise (FPN), and temporal noise (TN)—were characterized using the standard “S-curve” method [8]. This technique involves injecting a series of electrical pulses of increasing charge ($Q_{\rm injected}$) into each pixel and measuring the hit response. The injected charge is regulated by an external voltage pulse, $V_{\rm pulse}$, according to the relation $Q_{\rm injected}=V_{\rm pulse}\cdot C_{\rm test}$, where $C_{\rm test}$ is the injection capacitor with a designed value of 0.18 fF. The resulting S-shaped curve for each pixel is then analyzed to extract its specific threshold and noise characteristics. The final threshold value is converted to electrons via the value of $C_{\rm test}$.
2.2.1 Target Working-Point Tuning
@@
To thoroughly assess the impact of reverse bias on sensor performance, it is essential to tune the sensor to target working-points that exhibit a uniform average threshold across various substrate biases. Achieving such consistency poses a challenge due to the complex interplay of multiple parameters influencing the analog front-end. Therefore, the operating conditions—particularly the DAC settings that control the front-end bias condition—must be meticulously adjusted.
@@
This work focuses on the characterization of sectors 0-2. Sectors 0-2 share the common DAC configuration. The analog front-end design is identical for sectors 0-2, with variations existing only in their digital logic blocks for pixel functions and in their respective pixel pitches (see Tab. 2). Sector 3 is excluded from the present discussion; its front-end current was increased for radiation tolerance evaluations, leading to significantly higher power consumption.
@@
The M6 MOSFET, biased by VCLIP, clamps excessively high OUT_A signals (see Fig. 2). Given that the amplitude of OUT_A increases significantly with reverse bias due to a reduction in the sensor capacitance, VCLIP is set to 1.307 V. This is its maximum value and is chosen to eliminate the clamping effect of M6 on the OUT_A amplitude during characterization. To compensate for the increase of the M8 gate threshold induced by substrate reverse bias, VCASN2 is increased from its nominal 400 mV to 722 mV, thereby maintaining a sufficient gate voltage for the cascode structure of M8 and M9.
@@
Settings $V_{\rm sub}$ VCASN/mV ITHR/pA VCLIP/mV VCASN2/mV
@@ 0 V 400 500 1307 722
The threshold distributions for these target working-points are presented in Fig. 4. It is observed that these distributions at non-zero reverse biases are noticeably broader compared to zero bias, indicating increased pixel-to-pixel threshold variations. This broadening of the threshold distributions is quantified in Fig. 5, which plots the FPN as a function of bias voltage. The FPN is observed to increase with reverse bias, reflecting the growing pixel-to-pixel variations. The corresponding TN analysis is shown in Fig. 5. An increase in TN is observed under reverse bias conditions. The underlying reasons for this behavior will be discussed in Section 3.2 along with the change of the OUT_A signal amplitude.
@@
2.2.2 ITHR Scan
To investigate the sensor performance across a range of detection thresholds, the ITHR value was systematically scanned starting from each of these target working-points. This procedure yielded a series of operating conditions—7 distinct $V_{\rm sub}$ levels, each with 11 different ITHR settings—resulting in a set of working-points with varying thresholds.
Fig. 6 illustrates the impact of the ITHR value on the mean threshold, FPN and TN, respectively. Both FPN and TN generally exhibit an increasing trend with higher threshold values.
@@ 3 Experimental Results and Discussion
Reducing the sensor capacitance and expanding the depletion zone are the most intrinsic and direct consequences of applying a reverse bias on the substrate. After adjusting thresholds under different reverse biases and determining a set of working-points by the ITHR value scan described in Sec. 2.2.2, a thorough sensor characterization regarding the impact of the substrate reverse bias voltage was conducted, including leakage current, analog response, fake-hit rate and charge collection efficiency.
@@ 3.1 Leakage Current
The results, presented in Fig. 7, indicate that the leakage current from the pixel area is small (less than 3 fA per pixel) and increases with the reverse bias applied to the substrate. The current observed at substrate bias 0V is a result of the potential difference between the collection node, which is held at +1V by the VReset pin, and the grounded substrate. This observed increase with reverse bias is consistent with the understanding that a primary component of leakage current originates from the thermal generation of electron-hole pairs within the depletion region of the pixel diode. As the substrate reverse bias increases, the depletion region expands, thereby enlarging the volume for carrier generation and consequently leading to a higher measured leakage current [10].